A compound Q was found to be composed of 4.8g of element A and 6.4g of element B. What is the percentage composition of compound Q


1
Expert's answer
2021-11-09T10:24:58-0500

Total mass = 4.8 + 6.4 = 11.2 g

Proportion:

11.2 g – 100 %

4.8 g – x

"x = \frac{4.8 \times 100}{11.2} = 42.85 \;%"

Element A: 42.85 %

Element B: 100 – 42.85 = 57.15 %